A polygon has n sides. Two of its exterior angles are 72 and 35 while the remaining exterior angles are each equal to 23. Find the value of n.
Answer by Alan3354(30993)

A polygon has n sides. Two of its exterior angles are 72 and 35 while the remaining exterior angles are each equal to 23. Find the value of n.
-------------
The sum of exterior angles is 360 degs
---- 360 - (72+35) = 253 degs
253/23 = 11 angles of 23 degs
11 + 2 = 13 sides
n = 13